# InterDyn: Controllable Interactive Dynamics with Video Diffusion Models

### v1

@@ Abstract
Predicting the dynamics of interacting objects is essential for both humans and intelligent systems. However, existing approaches are limited to simplified, toy settings and lack generalizability to complex, real-world environments. Recent advances in generative models have enabled the prediction of state transitions based on interventions, but focus on generating a single future state which neglects the continuous motion and subsequent dynamics resulting from the interaction. To address this gap, we propose InterDyn, a novel framework that generates videos of interactive dynamics given an initial frame and a control signal encoding the motion of a driving object or actor. Our key insight is that large video foundation models can act as both neural renderers and implicit physics “simulators” by learning interactive dynamics from large-scale video data. To effectively harness this capability, we introduce an interactive control mechanism that conditions the video generation process on the motion of the driving entity. Qualitative results demonstrate that InterDyn generates plausible, temporally consistent videos of complex object interactions while generalizing to unseen objects. Quantitative evaluations show that InterDyn outperforms baselines that focus on static state transitions. This work highlights the potential of leveraging video generative models as implicit physics engines. Code and trained models will be released at: https://interdyn.is.tue.mpg.de/.
Figure 1: InterDyn. Given an image and a “driving motion” in the form of a mask sequence, our model generates scene dynamics that are physically plausible without 3D reconstruction and physics simulation. We investigate the generated “interactive dynamics” under simulated settings and in-the-wild HOI scenarios.
**footnotetext: Equal contribution
@@ 1 Introduction
Humans have the remarkable ability to intuitively predict the future dynamics of observed systems. With just a single image, we can anticipate and imagine how objects will move over time – not only their motion but also their interactions with the environment and other elements in the scene. Inferring this requires an advanced form of scene-level reasoning beyond merely recognizing the semantics and geometry of static elements; it involves a deep physical and causal understanding of how each object will interact given the environment, object properties, and forces.
@@
There has been a growing interest in developing machine learning systems that emulate similar levels of dynamic understanding given visual observations, such as images or videos. Early work [81] addressed this by first reconstructing a 3D representation from the image, then predicting future states with a physics simulator and finally generating the output video with a rendering engine. This relies heavily on explicit reconstruction and simulation, which is computationally intensive, prone to errors in the reconstruction, and may not generalize well. More recent methods [43, 2, 22, 46, 36] leverage keypoint or latent representations within graph relational frameworks; however, they have only been trained and validated in over-simplified, synthetic environments, showing limited generalizability to complex real-world scenarios.
@@
Instead, the advent of powerful generative models [4, 15, 50, 62, 1] opens new avenues for synthesizing interactions under complex scenarios. For example, Sudhakar et al. [66] recently proposed CosHand, a controllable image-to-image model based on Stable Diffusion [62] that infers state transitions of an object. The task here is defined as follows: given an image of a hand interacting with an object, alongside a hand mask of the current frame and a mask of the hand at a future frame, generate a modified input image that satisfies the mask, with realistic interactions. The challenge, as in early intuitive physics works, lies in accurately modeling how the objects will change after forces are applied. However, we argue that static state transitions are insufficient for this task, as they fail to capture the continuous dynamic processes inherent to the problem –see e.g. Fig. 2. Investigating interactive dynamics within a two-state setting is highly limiting, since dynamics can extend beyond the period of direct contact –for example, predicting the motion occurring while a person pours water requires a physical understanding that goes beyond the state of the hand at a future frame. The driving force, in this case the hand, may interact with the system only briefly, but the system’s subsequent dynamics continue according to physical laws and may even influence other parts via force propagation.
@@
In this paper, we explore controllable synthesis of interactive dynamics; generating a video from an input image and a dynamic control signal (e.g., a moving hand mask) to model realistic object dynamics. In particular, we propose InterDyn, a novel framework for synthesizing controllable dynamic interactions that leverages the physical and dynamics “knowledge” of a large video foundation model [4]. Unlike prior approaches that rely on explicit physical simulation [81] or are constrained to static state transitions [66], we leverage video foundation models to generate dynamic processes implicitly. Specifically, we extend Stable Video Diffusion (SVD[4] with a dynamic control branch and fine-tune it on diverse scenes, enabling the synthesis of complex interactions aligned with the control signal.
@@
We start our investigation by fine-tuning InterDyn on a simple synthetic scenario of cubes, cylinders, and spheres: the CLEVRER dataset [94]. To control the motion we add a mask driving signal that manipulates the movement of some (but not all) of the objects in the scene. We then evaluate how the synthesized trajectories of uncontrolled objects change under various interactions, including multiple objects colliding with each other. This multi-object collision setting allows us to “probe” the physical understanding and causal effects of the video diffusion model, and our qualitative experiments have shown InterDyn’s ability for counterfactual future prediction and physical force propagation.
@@
Further, we evaluate how the system performs in a highly complex, real-world scenario, such as Human-Object Interaction (HOI). Here, the dexterity of hand motions and the diversity of objects vastly increase the complexity of the problem. We fine-tune the model on a commonly used HOI video dataset [20] and compare it with the state-of-the-art baseline CosHand [66] as well as other recent baselines. In both scenarios, we quantify our investigations using standard image and video metrics, as well as a motion fidelity metric based on point tracking. InterDyn surpasses the previous SOTA over 37.5% on LPIPS and 77% on FVD on the Something-Something-v2 dataset [20]. Our experiments also demonstrate diverse physical plausible generations of interactive dynamics, probing into SVD’s “understanding” of physics and dynamics.
@@
In summary, we present InterDyn, the first framework that employs video generative models to simulate object dynamics without explicit 3D reconstruction or physical simulation. We demonstrate how the inherent ”knowledge” within video foundation models can be leveraged to predict complex object interactions and movements over time, implicitly modeling physical and causal dynamics. We perform comprehensive experiments on multi-object collision datasets and hand-object manipulation datasets, demonstrating the effectiveness of our approach.
@@ 2 Related Work
Modeling human-object interactions (HOI). Human-object interactions have been widely studied within the context of 3D reconstruction [25, 26, 69, 91, 16, 17, 85], where the goal is to recover realistic geometry of hands and objects. The field of 3D HOI synthesis has also received increasing attention, including the generation of static [67, 37, 104, 41] or dynamic [103, 101, 68, 59] hand poses conditioned on 3D objects, whole-body interactions [87], or more recently, hand-object meshes given textual descriptions [83, 9, 93, 53, 14]. Few works address HOI synthesis in the 2D domain. GANHand [13] predicts 3D hand shape and pose given an RGB image of an object, while AffordanceDiffusion [92] estimates a 2D hand instead, by leveraging a diffusion model. Kulal et al. [44] take as input an image of a human and a scene separately and generate a composite image that positions the human with correct affordances. Also relevant is HOIDiffusion [99], in which a texture-less rendering of a 3D hand and object is converted to a realistic image using a text description. Most closely related to us is CosHand [66], which takes as input an RGB image of a hand-object interaction, hand mask at the current state, and hand mask of the future state, and generates an RGB image of the future state. Unlike us, they cannot generate post-interaction object dynamics and are more heavily dependent on hand mask quality. Importantly, none of these works study dynamics, generating instead discrete state transitions that fail to capture the nuanced, temporally coherent behaviors observed in interactions.
@@
Controllable video generation. Video generation has advanced significantly in recent years, with diffusion models leading to substantial improvements in unconditional [30, 96], text-based [64, 29, 4, 34, 11, 102, 5, 1, 19, 89, 73, 24, 82] and image-based [4, 86, 75, 1, 19, 23] generation. These advances have raised the question of how to incorporate more nuanced control into video generation. Some text-to-video approaches are trained by “inflating” text-to-image (T2V) models [82, 35, 12, 23, 7, 24], and can thus be integrated with conditional T2V models such as ControlNet [98] or T2V-Adapter [56]. Control can also be achieved by conditioning on trajectories [84, 54, 95] or bounding-boxes [74], fine-tuning on appropriate datasets. VideoComposer [75] incorporates multiple condition types, including text, depth, style, and temporal conditions via motion vectors. Camera motion control has also been explored, with AnimateDiff [76] employing LoRA [31] modules to control camera movement, while MotionCtrl [77] and CameraCtrl [27] directly embed the camera information for more precise control. Additionally, several works target human animation from a pose control signal, such as DreamPose [39], MagicPose [88], and AnimateAnyone [32]; however, they do not account for interactions.
@@ 3 Controllable Interactive Dynamics
In this work, we extend Stable Video Diffusion [4] (SVD) to enable controllable interactive dynamics and explore the versatility of these models across a range of scenarios. SVD is a publicly available latent diffusion model [62] that extends Stable Diffusion 2.1 to video generation by interleaving the network with temporal layers. Given a static input image of a scene as conditioning, SVD denoises a sequence of $T$ frames $\boldsymbol{y}\in\mathbb{R}^{T\times H\times W\times 3}$ to generate a video that follows the initial frame. The conditioning image is fed into the denoising U-Net by concatenating its latent to each of the frames in the noised input, and by supplying its CLIP [60] embedding to the U-Net’s cross-attention layers. In addition, SVD is conditioned on the FPS and motion ID of the video, where the motion ID represents the amount of motion present in the sequence.
@@
Given an input image, $\boldsymbol{x}\in\mathbb{R}^{H\times W\times 3}$, and a driving motion in the form of a pixel-wise corresponding control signal $\boldsymbol{c}\in\mathbb{R}^{T\times H\times W}$, we task InterDyn with generating a video sequence, $\boldsymbol{y}\in\mathbb{R}^{T\times H\times W\times 3}$, depicting plausible object dynamics. Through this task, we aim to learn the conditional distribution between a driving motion, such as that of a human hand, and the consequent motion of manipulated objects. In our task, the model needs to synthesize plausible object movement and appearance without any indication other than the driving motion, while maintaining physical and visual consistency with the input image.
@@
InterDyn extends SVD with an additional control signal $\boldsymbol{c}\in\mathbb{R}^{T\times H\times W\times 3}$ by integrating a ControlNet-like branch [97]. The SVD weights remain frozen to preserve its learned dynamics prior. Following [97], we introduce a trainable copy of the SVD network connected to the original one via zero-convolutions. We use a small CNN, $\mathcal{E}(\cdot)$, to encode the control signal $\boldsymbol{c}$ into the latent space, which is then added to the noisy input latent that is passed to the ControlNet encoder. Importantly, similar to SVD, the control branch interleaves convolutional, spatial, and temporal blocks, enabling InterDyn to be temporally-aware when processing the control signal. This allows InterDyn to be robust to noisy control signals, see Figure 6.
We set the motion ID to a constant $40$ during training and inference. To facilitate classifier-free guidance [28], we randomly drop the input image with a probability of 5%. At inference, we start from Gaussian noise and use InterDyn to generate a video with plausible object dynamics. We choose masks as conditioning signals due to their accessibility. However, our method can potentially be extended to incorporate diverse types of signals. Our experiments demonstrate that the choice of conditioning signal does not significantly impact performance, see Appendix B.
@@ 4 Experiments
The primary goal of this work is to synthesize scene-level interactive dynamics by leveraging the implicit physical understanding of a pre-trained video generative model. We begin by probing the model’s ability to predict physically plausible outcomes within simulated environments, specifically using the CLEVRER dataset [94]. We test counterfactual future predictions based on different interactions and generate dynamics resulting from force propagation. Motivated by these promising results, we extend our investigation to complex, real-world hand-object interaction scenarios using the Something-Something-v2 dataset [20], conducting comprehensive comparisons with existing baselines that pursue similar objectives. Additionally, we showcase diverse physical examples to demonstrate the capabilities of InterDyn in generating realistic interactive dynamics.
4.1 Training details
We initialize InterDyn with the 14-frame publicly available [72] image-to-video weights of SVD [4]. We use the Adam optimizer [42] with a learning rate of $1\times 10^{-5}$. For all experiments, we use the EDM framework [40] with a noise distribution defined by $\log\sigma\sim\mathcal{N}(0.7,1.6^{2})$. We train on two NVIDIA H100 GPUs with a per-GPU batch size of 4. Dynamic events vary in duration; with brief actions such as dropping an object spanning just a couple of frames and longer actions such as moving an object spanning several seconds. Here, the context length and FPS together define the temporal resolution of dynamics. To strike a balance between capturing short- and long-range events, we subsample videos to 7 FPS for all datasets.
(a) Counterfactual Dynamics. Top: ground-truth collision sequence. Bottom: two different futures generated by InterDyn(1) w/ brown-rubber-cylinder and (2) w/o brown-rubber-cylinder. Note how the trajectory (shown as a red line) of the uncontrolled red-metal-sphere changes.
(b) Force Propagation. Top: ground-truth collision sequence. Bottom: InterDyn can generate plausible force propagation to and amongst uncontrolled objects (cyan-rubber-cylinder, pink-metal-cylinder, and pink-rubber-cube).

### v3

@@ Abstract
Predicting the dynamics of interacting objects is essential for both humans and intelligent systems. However, existing approaches are limited to simplified, toy settings and lack generalizability to complex, real-world environments. Recent advances in generative models have enabled the prediction of state transitions based on interventions, but focus on generating a single future state which neglects the continuous dynamics resulting from the interaction. To address this gap, we propose InterDyn, a novel framework that generates videos of interactive dynamics given an initial frame and a control signal encoding the motion of a driving object or actor. Our key insight is that large video generation models can act as both neural renderers and implicit physics “simulators”, having learned interactive dynamics from large-scale video data. To effectively harness this capability, we introduce an interactive control mechanism that conditions the video generation process on the motion of the driving entity. Qualitative results demonstrate that InterDyn generates plausible, temporally consistent videos of complex object interactions while generalizing to unseen objects. Quantitative evaluations show that InterDyn outperforms baselines that focus on static state transitions. This work highlights the potential of leveraging video generative models as implicit physics engines. Project page: https://interdyn.is.tue.mpg.de/.
Figure 1: We present InterDyn, a framework for synthesizing realistic interactive dynamics without 3D reconstruction and physics simulation. Our core principle is to rely on the implicit physics knowledge embedded in large-scale video generative models. Given an image and a “driving motion”, our model generates the consequential scene dynamics. We investigate the generated interactive dynamics in a simple object collision scenario (bottom) and complex in-the-wild human-object interaction (top).
**footnotetext: Equal contribution\$$\dagger$\$$\dagger$\$footnotetext: Project lead
@@ 1 Introduction
Humans have the remarkable ability to predict the future dynamics of observed systems intuitively. With just a single image, we can anticipate and imagine how objects will move over time – not only their motion but also their interactions with the environment and other elements in the scene. Inferring this requires an advanced form of scene-level reasoning beyond merely recognizing the semantics and geometry of static elements; it involves a deep physical and causal understanding of how each object will interact given the environment, object properties, and forces.
@@
There has been a growing interest in developing machine learning systems that emulate similar levels of dynamic understanding given visual observations, such as images or videos. Early work [84] addressed this by first reconstructing a 3D representation from the image, then predicting future states with a physics simulator and finally generating the output video with a rendering engine. This relies heavily on explicit reconstruction and simulation, which is computationally intensive, prone to errors, and may not generalize well. More recent methods [46, 2, 24, 49, 39] leverage keypoint or latent representations within graph relational frameworks; however, they have only been trained and validated in over-simplified, synthetic environments, showing limited generalizability to complex real-world scenarios.
@@
Instead, the advent of powerful generative models [5, 16, 54, 66, 1] opens new avenues for synthesizing interactions under complex scenarios. For example, Sudhakar et al. [70] recently proposed CosHand, a controllable image-to-image model based on Stable Diffusion [66] that infers state transitions of an object. The task here is defined as follows: given an image of a hand interacting with an object, alongside a hand mask of the current frame and a mask of the hand at a future frame, generate a modified input image that satisfies the mask, with realistic interactions. The challenge, as in early intuitive physics works, lies in accurately modeling how the objects will change after forces are applied. However, we argue that static state transitions are insufficient for this task, as they fail to capture the continuous dynamic processes inherent to the problem, e.g. see Fig. 2. Investigating interactive dynamics within a two-state setting is highly limiting, since dynamics can extend beyond the period of direct contact – for example, predicting the motion occurring while a person pours water requires a physical understanding that goes beyond the state of the hand at a future frame. The driving force, in this case the hand, may interact with the system only briefly, but the system’s subsequent dynamics continue according to physical laws and may even influence other parts via force propagation.
@@
In this paper, we explore controllable synthesis of interactive dynamics–generating a video from an input image and a dynamic control signal (e.g. a moving hand mask) to model realistic object dynamics. In particular, we propose InterDyn, a novel framework for synthesizing controllable dynamic interactions that leverages the physical and dynamics “knowledge” of a large video generation model [5]. Unlike prior approaches that rely on explicit physical simulation [84] or are constrained to static state transitions [70], we leverage video generation models to generate dynamic processes implicitly, see Fig. 1. Specifically, we extend Stable Video Diffusion (SVD[5] with a dynamic control branch and fine-tune it on diverse scenes, enabling synthesis of complex interactions aligned with the control signal.
@@
We start our investigation by fine-tuning InterDyn on a simple synthetic scenario of cubes, cylinders, and spheres: the CLEVRER dataset [97]. To control the motion we add a mask driving signal that manipulates the movement of some (but not all) of the objects in the scene. We then evaluate how the synthesized trajectories of uncontrolled objects change under various interactions, including multiple objects colliding with each other. This multi-object collision setting allows us to “probe” the physical understanding and causal effects of the video diffusion model, and our qualitative experiments show InterDyn’s ability for counterfactual future prediction and physical force propagation.
@@
Next, we evaluate how the system performs in a difficult real-world scenario, such as Human-Object Interaction (HOI). Here, the dexterity of hand motions and the diversity of objects vastly increase the complexity of the problem. We fine-tune the model on a commonly used HOI video dataset [21] and compare with the state-of-the-art baseline CosHand [70], as well as two text-control based interactive dynamics generation methods: Seer [23] and DynamiCrafter [89]. We quantify our investigations using standard image and video metrics, as well as a motion fidelity metric based on point tracking. InterDyn surpasses the previous SOTA over 37.5% on LPIPS and 77% on FVD on the Something-Something-v2 (SSV2) dataset [21]. Our experiments also demonstrate diverse and physically plausible generations of interactive dynamics, probing into SVD’s “understanding” of physics and dynamics.
@@
In summary, we present InterDyn, a framework that employs video generative models to simulate object dynamics without explicit 3D reconstruction or physical simulation. We demonstrate how the inherent “knowledge” within video generation models can be leveraged to predict complex object interactions and movements over time, implicitly modeling physical and causal dynamics. We perform comprehensive experiments on multi-object collision datasets and hand-object manipulation datasets, demonstrating the effectiveness of our approach.
@@ 2 Related Work
Modeling human-object interactions (HOI). Human-object interaction has been widely studied within the context of 3D reconstruction [27, 28, 73, 94, 17, 18, 88], where the goal is to recover realistic geometry of hands and objects. The field of 3D HOI synthesis has also received increasing attention, including the generation of static [71, 40, 106, 44] or dynamic [105, 103, 72, 63] hand poses conditioned on 3D objects, whole-body interactions [90], or more recently, hand-object meshes given textual descriptions [86, 10, 96, 57, 15]. Few works address HOI synthesis in the 2D domain. GANHand [14] predicts 3D hand shape and pose given an RGB image of an object, while AffordanceDiffusion [95] estimates a 2D hand using a diffusion model. Kulal et al. [47] take as input an image of a human and a scene separately and generate a composite image that positions the human with correct affordances. Also relevant is HOIDiffusion [101], in which a texture-less rendering of a 3D hand and object is converted to a realistic image using a text description. Most closely related to us is CosHand [70], which takes as input an RGB image of a hand-object interaction, a hand mask at the current state, and the hand mask of the future state, and generates an RGB image of the future state. Unlike us, they cannot generate post-interaction object dynamics. Importantly, none of these works study dynamics, generating instead discrete state transitions that fail to capture the nuanced, temporally coherent behaviors observed in interactions.
@@
Controllable video generation. Video generation has advanced significantly in recent years, with diffusion models leading to substantial improvements in unconditional [33, 99], text-based [68, 32, 5, 37, 12, 104, 6, 23, 1, 20, 92, 76, 26, 85, 89] and image-based [5, 89, 78, 1, 20, 25] generation. These advances have raised the question of how to incorporate more nuanced control into video generation. Some text-to-video approaches are trained by “inflating” text-to-image (T2V) models [85, 38, 13, 25, 8, 26], and can thus be integrated with conditional T2V models such as ControlNet [100] or T2V-Adapter [60]. Control can also be achieved by conditioning on trajectories [87, 58, 98] or bounding-boxes [77], fine-tuning on appropriate datasets. VideoComposer [78] incorporates multiple condition types, including text, depth, style, and temporal conditions via motion vectors. Camera motion control has also been explored, with AnimateDiff [79] employing LoRA [34] modules to control camera movement, while MotionCtrl [80] and CameraCtrl [29] directly embed the camera information for more precise control. Additionally, several works target human animation from a pose control signal, such as DreamPose [42], MagicPose [91], and AnimateAnyone [35], but do not account for interactions.
@@ 3 Controllable Interactive Dynamics
Task. Given an input image, $\boldsymbol{x}\in\mathbb{R}^{1\times H\times W\times 3}$, and a driving motion in the form of a pixel-wise corresponding control signal $\boldsymbol{c}\in\mathbb{R}^{N\times H\times W\times 3}$, we task InterDyn with generating a video sequence, $\boldsymbol{y}\in\mathbb{R}^{N\times H\times W\times 3}$, depicting plausible object dynamics. Through this task, we aim to learn the conditional distribution between a driving motion, such as that of a human hand, and the consequent motion of manipulated objects. In other words, the model needs to synthesize plausible object movement and appearance without any indication other than the driving motion, while maintaining physical and visual consistency with the input image.
Stable Video Diffusion. We extend Stable Video Diffusion [5] (SVD) to enable controllable interactive dynamics and explore the versatility of this model across a range of scenarios. SVD is a publicly available U-Net-based latent diffusion model [66] that extends Stable Diffusion 2.1 to video generation by interleaving the network with temporal layers. Given a static input image of a scene, SVD denoises a sequence of $N$ frames $\boldsymbol{y}\in\mathbb{R}^{N\times H\times W\times 3}$ to generate a video that follows the initial frame. The input image is fed into the denoising U-Net by concatenating its latent to each of the frames in the noised input, and by supplying its CLIP [64] embedding to the U-Net’s cross-attention layers. In addition, SVD is conditioned on the video’s FPS and motion ID, where the motion ID represents the amount of motion in the video. We found a motion ID of 40 to align well with our frozen SVD prior.
@@
Control. InterDyn extends SVD with an additional control signal $\boldsymbol{c}\in\mathbb{R}^{N\times H\times W\times 3}$ by integrating a ControlNet-like branch [100]. An overview of our pipeline is presented in Fig. 3. The SVD weights remain frozen to preserve its learned dynamics prior. Following [100], we introduce a trainable copy of the SVD encoder $E$, connected to SVD’s frozen decoder via skip connections, and modulated by zero-initialized convolutions. We use a small CNN, $\mathcal{E}(\cdot)$, to encode the control signal $\boldsymbol{c}$ into the latent space, which is then added to the noisy input latent that is passed to the ControlNet encoder. Similar to SVD, the control branch interleaves convolutional, spatial, and temporal blocks, enabling InterDyn to process the control signal in a temporal-aware manner. This helps InterDyn to be robust to noisy control signals, see Fig. 6. We opt for binary masks as conditioning signal due to their accessibility. However, our method can be extended to incorporate diverse types of signals. We find that for hand-object interactions, the type of conditioning signal does not significantly impact performance, see Appendix A and Tab. S1.
Figure 3: Overview of InterDyn. Given an input image depicting a scene, such as a hand holding a remote, and a “driving motion,” such as a sequence of binary hand masks, InterDyn generates a video depicting plausible hand and object dynamics. Crucially, InterDyn receives no control signal for the object. Through this setup, we probe and assess the implicit knowledge of large video generation models on complex interactive dynamics. We use Stable Video Diffusion (SVD) as our frozen backbone and fine-tune a separate control signal encoder. Videos are iteratively denoised over $T$ timesteps, starting from Gaussian noise $\boldsymbol{\epsilon}\sim\mathcal{N}(0,I)$.
Inference. During inference, we start from an input image, control signal sequence, and randomly sampled Gaussian noise $\boldsymbol{\epsilon}\sim\mathcal{N}(0,I)$. Through iteratively applying InterDyn over $T$ denoising timesteps, we generate a video $\boldsymbol{y}$ depicting plausible hand and object dynamics, aligned with the control signal.
@@ 4 Experiments
The primary goal of this work is to synthesize scene-level interactive dynamics by leveraging the implicit physical understanding of a pre-trained video generative model. We begin by probing the model’s ability to predict physically plausible outcomes within simulated environments, specifically using the CLEVRER dataset [97]. We test force propagation amongst uncontrolled objects and examine counterfactual future prediction by generating videos for one input image with different control signals. Motivated by promising results, we extend our investigation to complex, real-world hand-object interaction scenarios using the Something-Something-v2 (SSV2) dataset [21], conducting comprehensive comparisons with existing baselines that pursue similar objectives. Additionally, we showcase diverse physical examples to demonstrate the capabilities of InterDyn in generating realistic interactive dynamics.
(a) Force Propagation. InterDyn generates force propagation to and amongst uncontrolled objects (cyan cylinder, pink cylinder, and pink cube).
(b) Counterfactual Dynamics. Two generations of InterDyn on the same input image: (1) w/ brown-cylinder control and (2) w/o brown-cylinder control.
@@
4.1 Implementation
We initialize InterDyn with the 14-frame image-to-video weights of SVD [5]. During training, we use the Adam optimizer [45] with a learning rate of $1\times 10^{-5}$. We use the EDM framework [43] with noise distribution $\log\sigma\sim\mathcal{N}(0.7,1.6^{2})$. We train on two 80GB H100 GPUs, with a per-GPU batch size of 4. Video length and FPS define the temporal resolution of dynamics; to balance short- and long-range events we subsample videos to 7 FPS. To facilitate classifier-free guidance [31], we randomly drop the input image with a probability of 5%. At inference, we apply the Euler scheduler [43] over 50 denoising timesteps.